OpenStudy (anonymous):

What is the slope of the line that passes through the points (–6,1) and (4,–4)? I tried to do it and got -5/-2 What did i do wrong? a.-2 b.2 c.1/2 d.-1/2

OpenStudy (aum):

Slope \(\Large m = \frac{y_2 - y_1}{x_2 - x_1} = \frac{-4 - 1}{4 - -6} = ?\)

OpenStudy (anonymous):

-5/10 Simplify -1/2 so it would be D.

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

Thank you :)

OpenStudy (aum):

You are welcome.